OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \tan (x-y) = \frac{y}{1+x^2}\]

OpenStudy (lgbasallote):

how to start this?

OpenStudy (anonymous):

tan(x - y) = y / (1 + x^2) differentiating implicitly sec^2(x - y)(1 - y ') = [ (1+x^2)y ' - 2xy ] /(1 + x^2)^2 sec^2(x - y) - y ' sec^2(x - y) = y ' / (1 + x^2) - 2xy / (1 + x^2)^2 y ' [ 1 / (1 + x^2) + sec^2(x - y) ] = 2xy / (1 + x^2)^2 + sec^2(x - y) multiply with (1 + x^2)^2 y ' [ 1+ x^2 + (1+x^2)^2 sec^2(x - y) ] = 2xy + (1+x^2)^2 sec^2(x - y)

OpenStudy (anonymous):

y ' = \[2xy + (1+x^2)^2 \sec^2(x - y)\over (1 + x^2) [ 1 + (1+x^2)\sec^2(x-y)]\]

OpenStudy (lgbasallote):

lol nice...

OpenStudy (anonymous):

Is it correct @lgbasallote

OpenStudy (lgbasallote):

yup i thiink it is

OpenStudy (lgbasallote):

it looked intimidating at first haha guess it's just long

OpenStudy (anonymous):

lol